# ModeTv2: GPU-accelerated Motion Decomposition Transformer for Pairwise Optimization in Medical Image Registration

### v1

@@ Abstract
Deformable image registration plays a crucial role in medical imaging, aiding in disease diagnosis and image-guided interventions. Traditional iterative methods are slow, while deep learning (DL) accelerates solutions but faces usability and precision challenges. This study introduces a pyramid network with the enhanced motion decomposition Transformer (ModeTv2) operator, showcasing superior pairwise optimization (PO) akin to traditional methods. We re-implement ModeT operator with CUDA extensions to enhance its computational efficiency. We further propose RegHead module which refines deformation fields, improves the realism of deformation and reduces parameters. By adopting the PO, the proposed network balances accuracy, efficiency, and generalizability. Extensive experiments on two public brain MRI datasets and one abdominal CT dataset demonstrate the network’s suitability for PO, providing a DL model with enhanced usability and interpretability. The code is publicly available.
@@ 1 Introduction
Deformable image registration (DIR) has always been an important focus in the society of medical imaging, which is widely used for the preoperative planning, disease diagnosis and intraoperative guidance (Sotiras et al., 2013; Ferrante and Paragios, 2017; Fu et al., 2020). The deformable registration is to solve the non-rigid deformation field to warp the moving image, so that the warped image can be anatomically similar to the fixed image, as shown in Fig. 1. By doing so, information from mono-/multi-modalities can be fused into the same coordinate system to assist doctors with diagnosis and treatment (Ungi et al., 2016; Yang et al., 2022; Song et al., 2022; Rao et al., 2022). Although currently there are many DIR methods available, registration accuracy and efficiency are still challenges.
@@
Traditional medical image registration methods (Bajcsy and Kovačič, 1989; Thirion, 1998; Rueckert et al., 1999; Beg et al., 2005; Rueckert et al., 2006; Avants et al., 2008; Vercauteren et al., 2009; Heinrich et al., 2016) consider registration as an iterative process, gradually achieving well-registered images by iteratively optimizing the similarity between a pair of images. We refer to this registration process as pairwise optimization (PO). Such conventional approaches often require a substantial number of iterations to obtain a reasonable deformation field, which is evidently time-consuming. However, they possess the advantage of direct applicability to different domains (i.e., different anatomical structures and image modalities), unlike current deep learning methods that require retraining on different domains.
Deep learning (DL) approaches treat registration as a learning task (Fu et al., 2020; Haskins et al., 2020), wherein these methods extensively train on a training set to obtain a registration model. Subsequently, upon obtaining the model, there is no need for iterative computing on a new pair of images; instead, the deformation field is directly predicted through one forward pass. Although DL methods largely accelerate the solving speed of deformation fields, there still remains challenges. Firstly, in comparison to traditional methods, the usability of DL methods are constrained to the specific scenario presented in their training domain. Re-training a network for a new domain entails extra data and time. Secondly, even for the same domain task, the accuracy and interpretability of existing DL methods could be further improved.
@@
To this end, on one hand, progressively optimizing the deformation field from coarse to fine is a commonly employed strategy in DL registration tasks (Mok and Chung, 2020; Kang et al., 2022; Lv et al., 2022; Chen et al., 2022b; Ma et al., 2023; Zheng et al., 2024; Wang et al., 2024). The underlying assumption is that the network can first address global deformation at low resolution and then handle local deformation at high resolution. Such assumption exhibits a strong inductive bias tailored to the registration task, a strategy also adopted by traditional methods. Models employing this strategy can converge to good results with fewer training epochs due to the strong inductive bias inherent in the coarse-to-fine optimization assumption. However, these methods neglect multiple motion patterns at low-resolution voxels, which may lead to cumulative registration errors (Zheng et al., 2024). Additionally, most current pyramid networks consume substantial computational resources.
On the other hand, with the popularity of Transformer (Dosovitskiy et al., 2021) structures, some recent registration studies have employed cross-attention mechanisms (Song et al., 2021; Shi et al., 2022; Khor et al., 2023; Zhu et al., 2023) to capture spatial relationships between moving and fixed image features, aiming to enhance the model’s inductive bias for registration. Specifically, in these methods, one set (either the moving or fixed image) is used as the Query set, while the other set provides the Key and Value sets. The correlation between the Query and Key sets is computed to weight the values in the Value set. This step, akin to finding one-to-one correspondences in registration, is recognized as a promising strategy. However, most current methods utilize above mechanism merely to enhance feature representation, with few considering its direct usage for solving the deformation field (Liu et al., 2022; Chen et al., 2022b).
In our previous work (Wang et al., 2023), we propose a novel motion decomposition Transformer (ModeT) for deformable image registration. This operator simultaneously considers the decomposition of motion patterns in features and explicitly calculates the deformation field from the correlations of features, showcasing interpretability. In this study, we further improve the registration method of our previous one (Wang et al., 2023) and explore its capability for pairwise optimization in image registration. Specifically, we re-implement ModeT operator with CUDA extensions to enhance its computational efficiency. Moreover, we design a RegHead module to refine the deformation field. The efficacy of the proposed method is evaluated on two public brain MRI datasets and one abdominal CT dataset. Our method provides a DL model with enhanced accuracy and usability. Our contributions can be summarized as follows:
@@ 2.1 Traditional Registration Methods
Traditional registration methods involve iteratively optimizing a constrained deformation model to maximize similarity between fixed and moving images. Various nonlinear deformation methods have been proposed, such as optical flow based Demons (Thirion, 1998), B-spline constrained free-form deformation (FFD) (Rueckert et al., 1999), large deformation diffeomorphic metric mapping (LDDMM) (Beg et al., 2005), and standard symmetric normalization (SyN) (Avants et al., 2008), etc. Commonly used similarity descriptors for measuring image similarity during registration include sum of squared differences (SSD) (Wolberg and Zokai, 2000), normalized mutual information (NMI) (Knops et al., 2006), normalized cross-correlation (NCC) (Yoo and Han, 2009), and modality independent neighborhood descriptor (MIND) (Heinrich et al., 2012).
The primary limitation of these methods lies in their lengthy processing time, especially when dealing with volumetric data. However, their advantage lies in the direct applicability to any pair of images without the need for pre-training, providing convenient to use. For instance, tools like the ANTs (Avants et al., 2009) library, which integrates SyN (Avants et al., 2008), and the Elastix (Klein et al., 2009) registration package, which integrates FFD registration (Rueckert et al., 1999), allow users to simply input the images to directly perform registration without concerning themselves with image anatomy or modality.
@@ 2.2.1 Single Stage
Single-stage DL registration networks are among the earliest to emerge. The most well-known single-stage network is VoxelMorph (Balakrishnan et al., 2019), which employs a Unet-like structure (Ronneberger et al., 2015) to regress the deformation field. This structure is easy to be implemented, but it is not good enough to model complex displacements.
@@ 2.2.2 Cascading and Recurrent Structure
To address the issue of large deformations, especially in the context of unsupervised registration tasks, some studies utilize cascading or recurrent neural networks. The recursive cascaded networks (RCN) (Zhao et al., 2019a) concatenates multiple networks to handle the large deformation problem. Later, the recurrent registration neural networks (R2N2) (Sandkühler et al., 2019) and the recurrent registration network (RRN) (Jiang and Veeraraghavan, 2022) use gated recurrent units (GRU) (Chung et al., 2014) and convolutional long short term memory (CLSTM) (Shi et al., 2015) modules, respectively, to construct recurrent networks for continuous deformation. The idea behind cascading and recurrent methods is to further optimize the registration results generated by the preceding networks. The advantage is to handle large deformation, but these methods are computationally demanding and involve redundant feature extraction.
@@ 2.2.3 Pyramid Structure
To alleviate the computational burden of cascading and recurrent methods, pyramid registration methods have been investigated. The pyramid registration aims to address large deformation at low resolution first and then handle local deformation at high resolution. Currently, there are two types of pyramid registration models. One is similar to cascading networks, but the input at each level is a pyramid of the image itself, solving the deformation field from coarse to fine, as described in (Mok and Chung, 2020; Hering et al., 2019). The other type constructs the deep feature pyramid and employs these features entirely for pyramid registration (Hu et al., 2019; Cao et al., 2021; Kang et al., 2022; Lv et al., 2022; Liu et al., 2022; Meng et al., 2022; Ma et al., 2023). These models usually consist of dual encoders and several flow estimators. Specifically, the deep feature maps of the moving and fixed images are first obtained, and these features serve as inputs to the flow estimators, progressively obtaining the deformation field from coarse to fine resolution. Additionally, some studies combine pyramid structures with recurrent networks, such as (Hu et al., 2022; Zhou et al., 2023). The main drawback of pyramid structures is that they are prone to accumulating registration errors, as each voxel at low resolution covers a large portion of the original image. If there is a registration error at low resolution, it will propagate level-by-level and may become unrecoverable at high resolution. Some studies design different components to address this issue, which will be discussed in the next subsection.
@@ 2.3.1 Convolution
Early registration methods utilize convolutional neural networks (CNNs) to construct models (Balakrishnan et al., 2019; Zhao et al., 2019a; Hu et al., 2019). With the further development of CNNs, some advanced visual modules have been continuously incorporated into registration models. Encouraged by optical flow estimation, Kang et al. (2022) introduce a correlation layer to enhance the correlation estimation of the registration model. Zheng et al. (2024) enhance the separability of the deformation field by adding dilated convolutions (Chen et al., 2018). However, the addition of dilated convolutions slows down the computation, especially at high-resolution layers. Some studies (Li et al., 2022) employ channel and spatial attention modules (Hu et al., 2018; Woo et al., 2018) to perform convolutional attention calculations to re-weight feature maps. Lv et al. (2022) design specific attention modules to refine feature maps and deformation fields, respectively.
@@ 2.3.2 Self-Attention
The vision Transformer (ViT) (Dosovitskiy et al., 2021) structure has been employed in registration models to consider long-range dependencies and capture richer features. Chen et al. (2021) replace the bottom layer of the Unet structure with ViT modules. Furthermore, Chen et al. (2022a) use the SwinEncoder (Liu et al., 2021) instead of the Unet encoder for feature enhancement. Compared to the ViT structure, Swin adopts a relative friendly computing method with a moving window, leading to a more reasonable computational load. Some studies (Li et al., 2022; Zhu and Lu, 2022) also adopt similar designs. Recently, Meng et al. (2023) and Ma et al. (2023) propose pyramid structures and employ SwinTransformer at each level in the decoder as a powerful flow field estimator.
@@ 2.3.3 Cross-Attention
Furthermore, some studies have begun to consider using cross-attention (Vaswani et al., 2017) to capture relationships between features from moving and fixed images. Several studies (Chen et al., 2023; Zheng et al., 2022; Chen et al., 2024) use cross-attention encoders to enhance the feature extraction. Song et al. (2021) employ two cross-attention modules at the bottom layer of the U-shaped structure to perform bidirectional attention calculations on feature maps from two modalities to capture cross-modal features. Shi et al. (2022) use cross-attention at each layer from the encoder to the decoder. Aforementioned methods essentially employ the attention mechanisms to enhance feature representation. Only few researches directly solve the deformation field through attention calculations; however, this is a strong inductive bias for the registration task. Liu et al. (2022) calculate the matching scores of the feature maps from moving and fixed images. The computed scores are then employed to re-weight the deformation grid for obtaining the final deformation field. Chen et al. (2022b) utilize the multiplication of the attention map and Value matrix in the Transformer to weight the predicted bases to generate the deformation field. However, its attention map calculation merely concatenates and projects the moving and fixed feature maps without employing a similarity computation module. In our previous study, we propose the ModeT (Wang et al., 2023) operator, which uses multi-head neighborhood attention to weight the deformation grid, obtaining multiple motion modes at each level. This allows direct calculation of the deformation field while addressing the issue of error propagation in registration. However, the ModeT does not consider neighborhood relationships when fusing multiple motion modes, and also it could be implemented in a more efficient way.
@@ 2.4 Generalization Studies in Deep Registration
Due to the nature of unsupervised learning, registration networks are allowed to continue unsupervised optimization on new data, aiming to bring registered images closer in terms of similarity (Balakrishnan et al., 2019). Ferrante et al. (2018) specifically investigate the pairwise optimization capability of a single-stage CNN network, with the goal of exploring whether DL model can conduct registration in a way similar to traditional methods. The results show continuous PO on the same domain can further improve accuracy. However, for registration models transferred from a completely different domain (different modality or organ), even when trained to convergence, they still fail to match the performance of models directly trained on the target domain. Subsequent studies (Fechter and Baltas, 2020; Mok et al., 2023) adopt similar PO strategies. However, these methods generally focus on addressing the domain shift issue and do not discuss the handling capacity for entirely different domains. Furthermore, advanced network designs such as attention mechanisms have not been considered.
@@ 3.1 Network Overview
The proposed deformable registration network is illustrated in Fig. 2. We employ a pyramidal registration structure, which has the advantage of reducing the scope of attention calculation required at each decoding level and therefore alleviating the computational consumption. Initially, two sets of feature maps, $F_{1},F_{2},F_{3},F_{4},F_{5}$, and $M_{1},M_{2},M_{3},M_{4},M_{5}$, are generated for the fixed image $I_{f}$ and moving image $I_{m}$, respectively. Subsequently, the ModeTv2 is utilized as the deformation field estimator for each level to obtain the residual deformation field $\varphi_{l}~{}(l=1,2,3,4,5)$ at each level. Specifically, as shown in Fig. 2(b), the feature maps $M_{5}$ and $F_{5}$ are input into the ModeTv2, resulting in the coarsest level residual deformation field, denoted as $\varphi_{1}$ , serving as the initial value for the total deformation field $\phi_{1}$. The deformed moving feature map $M_{4}$, using the upsampled $\hat{\phi}_{1}$, is input into the next-level ModeTv2 along with the fixed feature map $F_{4}$ to obtain $\varphi_{2}$. The composition of $\varphi_{2}$ with the previous upsampled total deformation field $\hat{\phi}_{1}$ yields the updated total deformation field $\phi_{2}$, employing the composition formula from (Zhao et al., 2019b). Similar operations are performed for feature maps $M_{3},M_{2},M_{1}$, and $F_{3},F_{2},F_{1}$. Ultimately, utilizing the obtained total deformation field $\phi$, the image $I_{m}$ is warped to produce the registered image.
To guide the network training, the normalized cross correlation $\mathcal{L}_{\text{ncc}}$ (Rao et al., 2014) and the deformation regularization $\mathcal{L}_{\text{reg}}$ (Balakrishnan et al., 2019) is used:

### v2

@@ Abstract
Deformable image registration plays a crucial role in medical imaging, aiding in disease diagnosis and image-guided interventions. Traditional iterative methods are slow, while deep learning (DL) accelerates solutions but faces usability and precision challenges. This study introduces a pyramid network with the enhanced motion decomposition Transformer (ModeTv2) operator, showcasing superior pairwise optimization (PO) akin to traditional methods. We re-implement ModeT operator with CUDA extensions to enhance its computational efficiency. We further propose RegHead module which refines deformation fields, improves the realism of deformation and reduces parameters. By adopting the PO, the proposed network balances accuracy, efficiency, and generalizability. Extensive experiments on three public brain MRI datasets and one abdominal CT dataset demonstrate the network’s suitability for PO, providing a DL model with enhanced usability and interpretability. The code is publicly available at https://github.com/ZAX130/ModeTv2.
@@ 1 Introduction
Deformable image registration (DIR) has always been an important focus in the society of medical imaging, which is widely used for the preoperative planning, disease diagnosis and intraoperative guidance [62, 21, 23]. The deformable registration is to solve the non-rigid deformation field to warp the moving image, so that the warped image can be anatomically similar to the fixed image, as shown in Fig. 1. By doing so, information from mono-/multi-modalities can be fused into the same coordinate system to assist doctors with diagnosis and treatment [65, 73, 60, 51]. Although currently there are many DIR methods available, registration accuracy and efficiency are still challenges.
@@
Traditional medical image registration methods [5, 64, 55, 7, 54, 2, 67, 27] consider registration as an iterative process, gradually achieving well-registered images by iteratively optimizing the similarity between a pair of images. We refer to this registration process as pairwise optimization (PO). Such conventional approaches often require a substantial number of iterations to obtain a reasonable deformation field, which is evidently time-consuming. However, they possess the advantage of direct applicability to different domains (i.e., different anatomical structures and image modalities), unlike current deep learning methods that require retraining on different domains.
Deep learning (DL) approaches treat registration as a learning task [23, 24], wherein these methods extensively train on a training set to obtain a registration model. Subsequently, upon obtaining the model, there is no need for iterative computing on a new pair of images; instead, the deformation field is directly predicted through one forward pass. Although DL methods largely accelerate the solving speed of deformation fields, there still remains challenges. Firstly, in comparison to traditional methods, the usability of DL methods are constrained to the specific scenario presented in their training domain. Re-training a network for a new domain entails extra data and time. Secondly, even for the same domain task, the accuracy and interpretability of existing DL methods could be further improved.
@@
To this end, on one hand, progressively optimizing the deformation field from coarse to fine is a commonly employed strategy in DL registration tasks [48, 35, 44, 12, 45, 77, 70]. The underlying assumption is that the network can first address global deformation at low resolution and then handle local deformation at high resolution. Such assumption exhibits a strong inductive bias tailored to the registration task, a strategy also adopted by traditional methods. Models employing this strategy can converge to good results with fewer training epochs due to the strong inductive bias inherent in the coarse-to-fine optimization assumption. However, most current pyramid networks consume substantial computational resources. Additionally, these DL methods neglect multiple motion patterns at low-resolution voxels, which may lead to cumulative registration errors [77]. In contrast, some traditional methods [50, 68, 32] have explored various types of motion characteristics.
On the other hand, with the popularity of Transformer [18] structures, some recent registration studies have employed cross-attention mechanisms [61, 58, 36, 80] to capture spatial relationships between moving and fixed image features, aiming to enhance the model’s inductive bias for registration. Specifically, in these methods, one set (either the moving or fixed image) is used as the Query set, while the other set provides the Key and Value sets. The correlation between the Query and Key sets is computed to weight the values in the Value set. This step, akin to finding one-to-one correspondences in registration, is recognized as a promising strategy. However, most current methods utilize above mechanism merely to enhance feature representation, with few considering its direct usage for solving the deformation field [42, 12].
In our previous work [69], we propose a novel motion decomposition Transformer (ModeT) for deformable image registration. This operator simultaneously considers the decomposition of motion patterns in features and explicitly calculates the deformation field from the correlations of features, showcasing interpretability. In this study, we further improve the registration method of our previous one [69] and explore its capability for pairwise optimization in image registration. Specifically, we re-implement ModeT operator with CUDA extensions to enhance its computational efficiency. Moreover, we design a RegHead module to refine the deformation field. The efficacy of the proposed method is evaluated on three public brain MRI datasets and one abdominal CT dataset. Our method provides a DL model with enhanced accuracy and usability. Our contributions can be summarized as follows:
@@ 2.1 Traditional Registration Methods
Traditional registration methods involve iteratively optimizing a constrained deformation model to maximize similarity between fixed and moving images. Various nonlinear deformation methods have been proposed, such as optical flow based Demons [64], B-spline constrained free-form deformation (FFD) [55], large deformation diffeomorphic metric mapping (LDDMM) [7], and standard symmetric normalization (SyN) [2], etc. Commonly used similarity descriptors for measuring image similarity during registration include sum of squared differences (SSD) [71], normalized mutual information (NMI) [40], normalized cross-correlation (NCC) [74], modality independent neighborhood descriptor (MIND) [25], and self-similarity context (SSC) [26].
The primary limitation of these methods lies in their lengthy processing time, especially when dealing with volumetric data. Although some methods attempt to improve the computational efficiency, the processing time can still be improved. For example, the study [26] leverages discretized optimization framework, coarse grid spacing, and quantization to accelerate processing, but still takes approximately 20 seconds to conduct multimodal registration. Note that the advantage of most traditional methods lies in the direct applicability to any pair of images without the need for pre-training, providing convenient to use. For instance, tools like the ANTs [3] library, which integrates SyN [2], and the Elastix [39] registration package, which integrates FFD registration [55], allow users to simply input the images to directly perform registration without concerning themselves with image anatomy or modality. Another advantage of some traditional registration methods is that they have explored various types of motion characteristics. [50] replace Gaussian smoothing with bilateral filtering, which considers both spatial smoothness and local intensity similarity to preserve discontinuities in the deformation field. [68] propose the use of isotropic total variation (TV) regularization to accurately capture non-smooth displacement fields. [32] incorporate discontinuities by enriching B-spline basis functions with additional degrees of freedom.
@@ 2.2.1 Single Stage
Single-stage DL registration networks are among the earliest to emerge. The most well-known single-stage network is VoxelMorph [6], which employs a Unet-like structure [53] to regress the deformation field. This structure is easy to be implemented, but it is not good enough to model complex displacements.
@@ 2.2.2 Cascading and Recurrent Structure
To address the issue of large deformations, especially in the context of unsupervised registration tasks, some studies utilize cascading or recurrent neural networks. The recursive cascaded networks (RCN) [75] concatenates multiple networks to handle the large deformation problem. Later, the recurrent registration neural networks (R2N2) [56] and the recurrent registration network (RRN) [34] use gated recurrent units (GRU) [15] and convolutional long short term memory (CLSTM) [59] modules, respectively, to construct recurrent networks for continuous deformation. The idea behind cascading and recurrent methods is to further optimize the registration results generated by the preceding networks. The advantage is to handle large deformation, but these methods are computationally demanding and involve redundant feature extraction.
@@ 2.2.3 Pyramid Structure
To alleviate the computational burden of cascading and recurrent methods, pyramid registration methods have been investigated. The pyramid registration aims to address large deformation at low resolution first and then handle local deformation at high resolution. Currently, there are two types of pyramid registration models. One is similar to cascading networks, but the input at each level is a pyramid of the image itself, solving the deformation field from coarse to fine, as described in [48, 28]. The other type constructs the deep feature pyramid and employs these features entirely for pyramid registration [31, 8, 35, 44, 42, 46, 45]. These models usually consist of dual encoders and several flow estimators. Specifically, the deep feature maps of the moving and fixed images are first obtained, and these features serve as inputs to the flow estimators, progressively obtaining the deformation field from coarse to fine resolution. Additionally, some studies combine pyramid structures with recurrent networks, such as [29, 79]. The main drawback of pyramid structures is that they are prone to accumulating registration errors, as each voxel at low resolution covers a large portion of the original image. If there is a registration error at low resolution, it will propagate level-by-level and may become unrecoverable at high resolution. Some studies design different components to address this issue, which will be discussed in the next subsection.
@@ 2.3.1 Convolution
Early registration methods utilize convolutional neural networks (CNNs) to construct models [6, 75, 31]. With the further development of CNNs, some advanced visual modules have been continuously incorporated into registration models. Encouraged by optical flow estimation, [35] introduce a correlation layer to enhance the correlation estimation of the registration model. [77] enhance the separability of the deformation field by adding dilated convolutions [13]. However, the addition of dilated convolutions slows down the computation, especially at high-resolution layers. Some studies [41] employ channel and spatial attention modules [30, 72] to perform convolutional attention calculations to re-weight feature maps. [44] design specific attention modules to refine feature maps and deformation fields, respectively.
@@ 2.3.2 Self-Attention
The vision Transformer (ViT) [18] structure has been employed in registration models to consider long-range dependencies and capture richer features. [10] replace the bottom layer of the Unet structure with ViT modules. Furthermore, [9] use the SwinEncoder [43] instead of the Unet encoder for feature enhancement. Compared to the ViT structure, Swin adopts a relative friendly computing method with a moving window, leading to a more reasonable computational load. Some studies [41, 81] also adopt similar designs. Recently, [47] and [45] propose pyramid structures and employ SwinTransformer at each level in the decoder as a powerful flow field estimator.
@@ 2.3.3 Cross-Attention
Furthermore, some studies have begun to consider using cross-attention [66] to capture relationships between features from moving and fixed images. Several studies [11, 78, 14] use cross-attention encoders to enhance the feature extraction. [61] employ two cross-attention modules at the bottom layer of the U-shaped structure to perform bidirectional attention calculations on feature maps from two modalities to capture cross-modal features. [58] use cross-attention at each layer from the encoder to the decoder. Aforementioned methods essentially employ the attention mechanisms to enhance feature representation. Only few researches directly solve the deformation field through attention calculations; however, this is a strong inductive bias for the registration task. [42] calculate the matching scores of the feature maps from moving and fixed images. The computed scores are then employed to re-weight the deformation grid for obtaining the final deformation field. [12] utilize the multiplication of the attention map and Value matrix in the Transformer to weight the predicted bases to generate the deformation field. However, its attention map calculation merely concatenates and projects the moving and fixed feature maps without employing a similarity computation module. In our previous study, we propose the ModeT [69] operator, which uses multi-head neighborhood attention to weight the deformation grid, obtaining multiple motion modes at each level. This allows direct calculation of the deformation field while addressing the issue of error propagation in registration. However, the ModeT does not consider neighborhood relationships when fusing multiple motion modes, and also it could be implemented in a more efficient way.
@@ 2.4 Generalization Studies in Deep Registration
Due to the nature of unsupervised learning, registration networks are allowed to continue unsupervised optimization on new data, aiming to bring registered images closer in terms of similarity [6]. [20] specifically investigate the pairwise optimization capability of a single-stage CNN network, with the goal of exploring whether DL model can conduct registration in a way similar to traditional methods. The results show continuous PO on the same domain can further improve accuracy. However, for registration models transferred from a completely different domain (different modality or organ), even when trained to convergence, they still fail to match the performance of models directly trained on the target domain. Subsequent studies [19, 49] adopt similar PO strategies. However, these methods generally focus on addressing the domain shift issue and do not discuss the handling capacity for entirely different domains. Furthermore, advanced network designs such as attention mechanisms have not been considered.
@@ 3.1 Network Overview
The proposed deformable registration network is illustrated in Fig. 2. We employ a pyramidal registration structure, which has the advantage of reducing the scope of attention calculation required at each decoding level and therefore alleviating the computational consumption. Initially, two sets of feature maps, $F_{1},F_{2},F_{3},F_{4},F_{5}$, and $M_{1},M_{2},M_{3},M_{4},M_{5}$, are generated for the fixed image $I_{f}$ and moving image $I_{m}$, respectively. Subsequently, the ModeTv2 is utilized as the deformation field estimator for each level to obtain the residual deformation field $\varphi_{l}~{}(l=1,2,3,4,5)$ at each level. Specifically, as shown in Fig. 2(b), the feature maps $M_{5}$ and $F_{5}$ are input into the ModeTv2, resulting in the coarsest level residual deformation field, denoted as $\varphi_{1}$ , serving as the initial value for the total deformation field $\phi_{1}$. The deformed moving feature map $M_{4}$, using the upsampled $\hat{\phi}_{1}$, is input into the next-level ModeTv2 along with the fixed feature map $F_{4}$ to obtain $\varphi_{2}$. The composition of $\varphi_{2}$ with the previous upsampled total deformation field $\hat{\phi}_{1}$ yields the updated total deformation field $\phi_{2}$, employing the composition formula from [76]. Similar operations are performed for feature maps $M_{3},M_{2},M_{1}$, and $F_{3},F_{2},F_{1}$. Ultimately, utilizing the obtained total deformation field $\phi$, the image $I_{m}$ is warped to produce the registered image.
To guide the network training, the normalized cross correlation $\mathcal{L}_{\text{ncc}}$ [52] and the deformation regularization $\mathcal{L}_{\text{reg}}$ [6] is used: